# Ola: Pushing the Frontiers of Omni-Modal Language Model with Progressive Modality Alignment

### v2

@@ 3.2 Progressive Omni-Modal Alignment
Rethinking Modality Gaps among Language, Vision and Audio. From our exploration, we recognize two critical issues in omni-modal training. 1) Modal balancing. As illustrated in Fig. 3, directly combining data from all modalities negatively affects benchmark performance. Therefore, we propose a rational training procedure that progressively equips the sense organ to the Ola model. We assert that texts and images are the core modalities in omni-modal learning, while speeches and videos are variants of texts and images, respectively. Learning to recognize texts and images ensures the model’s basic cross-modal ability, so we prioritize these harder cases. Subsequently, we gradually incorporate video, audio, and speech into the training for the omni-modal LLM. 2) Connections between audio and vision. Another problem lies in building connections between audio and vision, which has been overlooked by previous works. However, jointly learning audio and vision data can yield surprising results in omni-modal learning by providing a more comprehensive view across different modalities. For the Ola model, we consider video as the bridge between audio and vision, as videos contain natural, abundant, and highly relevant information between frames and the accompanying audio. We test our hypothesis by optimizing the training pipeline and preparing targeted training data, as introduced below.
@@
Stage 2: Continuous Training for Images and Videos. Based on a strong text-image multi-modal LLM, we continuously extend the capability for Ola with video data. We keep most of the experimental setting for supervised fine-tuning while freezing the vision encoder in this stage as the encoder is already fully trained beforehand. We mix the previous image data and the video data to preserve the original text-image performance.
Stage 3: Bridging Vision and Audio with Videos. The audio-relevant training is included in stage 3. We follow the training strategy for the visual MLP adapter while initializing the audio MLP adapter with a basic audio-speech recognition (ASR) task. Then we mix up the text & speech understanding, text & music understanding, audio & video joint comprehension, and the foremost text-image multi-modal tasks together for the formal training. Ola concentrates on learning audio recognition and identifying the relationships between vision and audio in this stage, resulting in a comprehensive image, video, and audio understanding model after training.
3.3 Data
The training data of Ola includes the general supervised fine-tuning data collected from open-source academic datasets in image, video, and audio categories. Additionally, we design a pipeline to generate cross-modal video-audio data for omni-modal alignment.
Image Data. We follow the simple setting in [38] for image MLP alignment. The MLP alignment data includes 800k image captioning pairs from the LAION dataset [62]. For the large-scale pre-training phase, we collect around 20M text-image data pairs from open-sourced and in-house data to build the basic capability of the model. For text-image supervised fine-tuning data, we collect abundant data from various tasks including captions, conversations, OCR, etc. The source of the training data involves the mixture of LLaVA-OneVision [32], Cauldron [31], Cambrian-1 [67], MAmmoTH-VL [26], PixMo [12], etc., resulting in 7.3M image training data in total.
Video Data. For text-video training data, we collect useful video datasets from LLaVA-Video-178K [84], VideoChatGPT-Plus [47], LLaVA-Hound [82], and Cinepile [60], with 1.9M video conversation pieces in total. We randomly sample 2/3 video-language data pairs from LLaVA-Video-178K, resulting in 1.2M high-quality training data, and we use the full set of other data sources. In stage 2 for multi-image training, we randomly sample 0.8M image data from stage 1 and mix it with the video datasets for continuous training to maintain the basic performance.
Audio Data. We prepare audio training data for comprehensive speech and music understanding. For text-speech understanding, we design multiple tasks including ASR from LibriSpeech [54] and GigaSpeech [5] datasets, audio captioning from AudioCaps [30] and Clotho [14] datasets, speech question answering from LibriSpeech [54] datasets, audio question answering from WavCaps [49] and AudioCaps [30] datasets. For text-music understanding, we collect tasks including music captioning from MusicCaps [1], music question answering from MillionSong [48] and MusicNet [66]. The overall audio training data involves 1.1M samples. The relevant text question-answering representations are collected from SALMONN [63].
@@ 3.3 Data
Generating Cross-Modal Video Data. Most existing video training data are annotated or synthesized solely from frame inputs, often overlooking the valuable information in accompanying audio. To address this, we designed a pipeline to generate cross-modal video data, aiming to uncover the intrinsic relationships between video and audio. This guides an omni-modal large language model in learning cross-modality information. Specifically, we developed two tasks for cross-modal learning: video-audio question answering and video speech recognition. We collected videos from the academic dataset LLaVA-Video-178k [84] and the open-ended video datasets from FineVideo [18]. Due to the lack of subtitles in the academic datasets, we used Whisper-v3 [58] to generate subtitles from the video audio and conducted a language-based cleaning procedure. We then employed a large language model to assess whether the subtitles were complete and informative. We gathered 41k pure videos from LLaVA-Video-178k, along with the original 42k videos from FineVideo. Subsequently, we used Qwen2-VL-72B [57] to generate questions and answers based on the video and corresponding subtitles. The model was instructed to focus on the subtitle inputs while using the videos as supplementary information. We created three question-answer pairs for each video, resulting in 243k cross-modal video-audio data points. Additionally, we included the original video subtitling tasks with 83k training data to help the model maintain its ASR ability in noisy environments. During training, the models processed multiple image frames, audio, and text inputs simultaneously, significantly enhancing their cross-modal capabilities.
The prepared 324k cross-modal video data is mixed with audio data implemented above for stage 3 training. We mix all the 1.1M pure text-audio training data and 324k cross-modal video-audio data for the comprehensive training stage. Additionally, we sample 400k image data from stage 1 to maintain the basic ability and create 200k image data with voice instructions to equip the model with interaction capability.
4 Experiments
@@ 4.2 Results on Omni Understanding
Benchmarks. We conduct extensive comparisons across image, video, and audio understanding benchmarks to demonstrate the omni-modal capabilities of the Ola model. For image benchmarks, we utilize comprehensive understanding datasets including MMBench-1.1 [40], MMMU [77], MMStar [6], MathVista [46], Hallusion Bench [25], AI2D [29], and OCRBench [41]. In the video domain, we evaluate using VideoMME [21], which involves multiple-choice questions on videos of varying lengths, and LongVideoBench [69] for assessing performance on extremely long video content, MVBench [34] for the general recognition ability. For audio benchmarks, we focus on two primary tasks relevant to audio LLMs. Librispeech [54] serves as a traditional audio-speech recognition (ASR) dataset, testing the model’s ability to accurately transcribe spoken language. AIR-Bench [74] provides a comprehensive evaluation of audio question-answering capabilities, incorporating speech, sound, and music inputs. The responses are evaluated with a GPT-based [52] scorer against ground truth answers. We report the mainstream evaluation metric on image and video benchmarks and report the mean metric in Tab. 2 for ASR and audio understanding tasks for simplicity.
Baselines. We selected a range of state-of-the-art multi-modal large language models across different modalities for comparison and reference. We categorized vision-language models into three groups: image-centric LLMs, video-centric LLMs, and comprehensive LLMs capable of handling both images and videos. For image understanding, we utilized Cambrian-1 [67] and Pixtral-12B [2]. For video understanding, VideoCCAM [20] and LLaVA-Video [84] are employed. Comprehensive models included LLaVA-OneVision [32], MiniCPM-V 2.6 [75], InternVL2.5 [9], and Qwen2.5-VL [65] which excel across various visual benchmarks. In the audio domain, we compared our work with state-of-the-art models such as SALMONN [63] and Qwen-2 Audio [11]. As an Omni-modal LLM, our model, Ola, was compared with state-of-the-art open-source omni-modal LLMs like Mini-Omni2 [72], VITA-1.5 [22], InternLM-XComposer2.5-OmniLive [81], which support image, audio, and text inputs. Additionally, LLaMA-Omni [72], an audio-text omni model, was noted for its strong speech generation capabilities.
Table 2: Analysis Results on Audio Benchmarks. We report the WER rate on test-clean, test-other, dev-clean, dev-other subsets of LibriSpeech, and the scores on AIR-Bench. In the table, ”--” indicates the model is capable of solving the tasks, while the result is lacking. ”✗” indicates that the model is not capable of the task.
@@
Results. We present the comprehensive results in Table 1, highlighting Ola’s competitive performance across major multi-modal benchmarks when compared to state-of-the-art specialist-modal LLMs. Specifically, in image benchmarks, Ola achieves 84.3% on MMBench-1.1 [40], 70.8% on MMStar [6], 57.0% on MMMU [77], 68.4% on MathVista [46], 86.1% on AI2D [29] and 827 on OCRBench [41], surpassing all the relative multi-modal LLMs in similar number of parameters. In video benchmarks, Ola attains an impressive 68.4% on VideoMME [21], showcasing its robust capability to handle both video and audio inputs simultaneously, and setting a new state-of-the-art performance among 7B models on the VideoMME benchmark. Ola also maintains a leading position compared to mainstream video LLMs including LLaVA-Video [84] and VideoCCAM [20] on LongVideoBench [69] and MVBench [34]. In audio benchmarks, Ola demonstrates strong audio-speech recognition and conversational abilities, with 3.1% mean WER rate on LibriSpeech [54] and 6.41 mean score on AIR-Bench [74], outperforming existing omni-modal LLMs, including LLaMA-Omni [17], which focuses on audio understanding. These results indicate a significant advancement over current omni-modal LLMs, underscoring the effectiveness of Ola’s training approach.
@@ 4.3 Analysis
Analysis on Audio Benchmarks. To demonstrate the effectiveness of our approach on audio and speech tasks, we conducted experiments using the LibriSpeech [54] and AIR-Bench [74] datasets, and we illustrate our results on Tab. 2. Specifically, we report the Word Error Rate (WER) on the test-clean, test-other, dev-clean, and dev-other subsets of LibriSpeech. Additionally, we present the GPT-4-eval scores on speech, sound, music, and mix sub-metrics in AIR-Bench. Our model, Ola, is compared against state-of-the-art audio models and omni-modal LLMs.
Notably, Ola demonstrates a significant advantage over existing omni-modal models, achieving a 1.9 WER on the test-clean subset of LibriSpeech and a 6.41 average score on AIR-Bench. This is in contrast to the previous state-of-the-art omni-modal models, which achieved a 2.5 WER and a 4.70 score, respectively. Ola’s performance is even approaching that of audio-specific models, highlighting its strong universality.
@@
Figure 5: Generative results on speech and visual understanding tasks. We illustrate results on speech and video understanding and show the strong ability of omni-modal Ola compared with conventional vision-language models.
@@
Showcases. We present the qualitative generation results of the Ola model on both speech and visual understanding tasks in Fig. 5. For speech understanding, we utilize sources from the AIR-Bench [74] benchmark, where Ola demonstrates precise speech recognition and effective emotional analysis, as well as reasoning about the questions posed. Ola tries to find out the answer for the instruction based on the audio inputs. In visual understanding tasks, we analyze an interview with a famous tennis player after the Olympic Games. We compare our approach with the state-of-the-art vision-language model [32]. The large vision-language models without audio inputs exhibit significant information loss due to the absence of audio input and recognition capabilities for audio. In contrast, Ola, with its omni-modal inputs, provides more accurate responses regarding nationality, context, and background from the speaker’s dialogue.
@@ 5 Conclusion
In this paper, we propose Ola, a comprehensive and powerful omni-modal language model that achieves competitive performance in image, video, and audio understanding tasks. Our solution, based on a progressive modality alignment strategy, offers a natural, efficient, and competitive pipeline for training omni-modal models. Enhancements in architectural design with omni-modal inputs and streaming decoding, along with high-quality cross-modal video data preparation, further extend Ola’s capabilities. We hope our work inspires future research on more general AI models.
@@ A.1 Model Details
The visual encoder for Ola is based on the SigLIP-400M [79] backbone and is further fine-tuned for native-resolution visual inputs. The patch size for the visual encoder is set to 16, the hidden dimension is 1152, and the MLP hidden dimension is 4304. The SigLIP-400M model consists of 27 transformer blocks and 16 attention heads. The audio encoder for Ola is built on the whisper-v3-large [58] model. The length of the input audio tensor for the whisper model is fixed at 480,000; therefore, we chunk the entire audio tensor into pieces and concatenate the audio features. The mel size for the whisper-v3-large model is set to 128, and the hidden dimension for the speech features after the whisper-v3-large model is 1280.
@@
We integrate Qwen-2.5-7B [64] into the Ola large language model. The Qwen-2.5-7B model has a hidden LLM dimension of 3,584 and an intermediate size of 18,944. It consists of 28 transformer layers. The basic architecture for the speech decoder mirrors that of the Qwen-2.5-7B LLM, but we use only a 2-layer transformer block for the speech decoder. During generation, the maximum number of classification categories for the unit speech tensor is set to 1,000. We use a pre-trained speech vocoder to convert the unit speech tensor into speech waveforms. During inference, speech outputs are generated whenever a punctuation mark is detected, ensuring that the features of streaming outputs are preserved.
@@ A.2 Training Details
As stated in the main paper, the progressive modality alignment procedure is conducted in four stages. The first image-text stage involves adapter pre-training and supervised fine-tuning. The adapter pre-training stage is conducted on 808k image captioning data collected from the LAION datasets. During pre-training, we unfreeze the parameters for the connector while keeping other parameters frozen. We set the training batch size to 256 and the overall learning rate to 1e-3. The supervised fine-tuning stage is conducted on 7.3 million image-text data pairs. During image-text training, input images are maintained at their original aspect ratio, with the maximum image size restricted to 1536×1536. We set the batch size to 128 and the overall learning rate to 2e-5. The stage 1 experiment is conducted on 64 NVIDIA A800 GPUs.

### v3

@@ 3.3 Bridging Modality Gaps for Omni-Modal Understanding
Rethinking Modality Gaps among Language, Vision and Audio. From our exploration, we recognize two critical issues in omni-modal training. 1) Connections between modalities. In conventional learning strategies for omni-modal models, the training samples emphasizing the connections between modalities are always lacking, especially in two major modalities: vision and audio. From our efforts, we observe that jointly learning audio and vision data can yield surprising results in omni-modal learning by providing a more comprehensive view across different modalities. For the Ola model, we consider video as the bridge between audio and vision, as videos contain natural, abundant, and highly relevant information between frames and the accompanying audio. We test our hypothesis by optimizing the training pipeline and preparing targeted training data, as introduced below. 2) Modal balancing. As illustrated in Fig. 1 (c), directly combining data from all modalities negatively affects benchmark performance. Therefore, we propose a rational training procedure that progressively equips the sense organ to the Ola model. We assert that texts and images are the core modalities in omni-modal learning, while speeches and videos are variants of texts and images, respectively. Learning to recognize texts and images ensures the model’s basic cross-modal ability, so we prioritize these harder cases. Subsequently, we gradually incorporate video, audio, and speech into the training for the omni-modal LLM.
3.3.1 Deriving Omni-Modal Model from Videos
Most existing video training data are annotated or synthesized solely from frame inputs, often overlooking the valuable information in accompanying audio. To address this, we designed a pipeline to generate cross-modal video data, aiming to uncover the intrinsic relationships between video and audio. This guides an omni-modal large language model in learning cross-modality information. Specifically, we developed two tasks for cross-modal learning: video-audio question answering and video speech recognition. We collected videos from the academic dataset LLaVA-Video-178k [80] and the open-ended video datasets from FineVideo [17]. Due to the lack of subtitles in the academic datasets, we used Whisper-v3 [55] to generate subtitles from the video audio and conducted a language-based cleaning procedure. We then employed a large language model to assess whether the subtitles were complete and informative. We gathered 41k pure videos from LLaVA-Video-178k, along with the original 42k videos from FineVideo. Subsequently, we used Qwen2-VL-72B [54] to generate questions and answers based on the video and corresponding subtitles. The model was instructed to focus on the subtitle inputs while using the videos as supplementary information. We created three question-answer pairs for each video, resulting in 243k cross-modal video-audio data. Additionally, we included the original video subtitling tasks with 83k training data to help the model maintain its ASR ability in noisy environments. During training, the models processed multiple frames, audio, and text inputs simultaneously, significantly enhancing their cross-modal capabilities.
3.3.2 Omni-Modal Training with Progressive Modality Alignment
We start from two fundamental and separated modalities, image and text, to build the basic knowledge for omni-modal models. Subsequently, we gradually expand the training sets to equip the model with an extended ability, including video frames that strengthen the visual understanding capability, speech data that connects the language and audio knowledge, and the video with audio that mixes up the information from language, video, and audio comprehensively.
@@ 3.3.2 Omni-Modal Training with Progressive Modality Alignment
Stage 2: Continuous Training for Images and Videos. Based on a strong text-image multi-modal LLM, we continuously extend the capability for Ola with video data. We keep most of the experimental setting for supervised fine-tuning while freezing the vision encoder in this stage as the encoder is already fully trained beforehand. We mix the previous image data and the video data to preserve the original text-image performance. We randomly sample 0.8M image data from stage 1 and mix it with the video datasets for continuous training to maintain the basic performance.
Stage 3: Bridging Vision and Audio with Videos. The audio-relevant training is included in stage 3. We follow the training strategy for the visual MLP adapter while initializing the audio MLP adapter with a basic audio-speech recognition (ASR) task. Then we mix up the text & speech understanding, text & music understanding, audio & video joint comprehension, and the foremost text-image multi-modal tasks together for the formal training. Ola concentrates on learning audio recognition and identifying the relationships between vision and audio in this stage, resulting in a comprehensive image, video, and audio understanding model. We mix up all the 324k cross-modal data, 1.1M pure text-audio data for the comprehensive training stage. Additionally, we sample 400k image data from stage 1 to maintain the basic ability and create 200k image data with voice instructions to equip the model with interaction capability.
4 Experiments
@@ 4.2 Results on Omni Understanding
Benchmarks. We conduct extensive comparisons across image, video, and audio understanding benchmarks to demonstrate the omni-modal capabilities of the Ola model. For image benchmarks, we utilize comprehensive understanding datasets including MMBench-1.1 [37], MMMU [73], MMStar [6], MathVista [43], Hallusion Bench [24], AI2D [28], and OCRBench [38]. In the video domain, we evaluate using VideoMME [20], which involves multiple-choice questions on videos of varying lengths, and LongVideoBench [66] for assessing performance on extremely long video content, MVBench [32] for the general recognition ability. For audio benchmarks, we focus on two primary tasks relevant to audio LLMs. Librispeech [51] serves as a traditional audio-speech recognition (ASR) dataset, testing the model’s ability to accurately transcribe spoken language. AIR-Bench [70] provides a comprehensive evaluation of audio question-answering capabilities, incorporating speech, sound, and music inputs. The responses are evaluated with a GPT-based [49] scorer against ground truth answers. We report the mainstream evaluation metric on image and video benchmarks and report the mean metric for audio understanding tasks for simplicity.
Baselines. We selecte a range of state-of-the-art MLLMs across different modalities for comparison and reference. We categorize vision-language models into three groups: image-centric LLMs, video-centric LLMs, and comprehensive LLMs capable of handling both images and videos. For image understanding, we utilized Cambrian-1 [65] and Pixtral-12B [2]. For video understanding, VideoCCAM [19] and LLaVA-Video [80] are employed. Comprehensive models included LLaVA-OneVision [31], MiniCPM-V 2.6 [71], InternVL2.5 [9], and Qwen2.5-VL [63] which excel across various visual benchmarks. In the audio domain, we compared our work with SOTA models such as SALMONN [61] and Qwen-2 Audio [11]. As an Omni-modal LLM, our model, Ola, was compared with SOTA open-source omni-modal LLMs like Mini-Omni2 [68], VITA-1.5 [21], InternLM-XComposer2.5-OmniLive [77], Qwen2.5-Omni [69]. Additionally, LLaMA-Omni [68], an audio-text omni model, was noted for its strong speech generation capabilities.
Table 2: Analysis Results on Audio Benchmarks. We report the WER rate on test-clean, test-other, dev-clean, dev-other subsets of LibriSpeech, the scores on AIR-Bench, MMAU, and AIR-Foundation.. In the table, ”--” indicates the model is capable of solving the tasks, while the result is lacking. ”✗” indicates that the model is not capable of the task.
@@
Results. We present the results in Table 1, highlighting Ola’s competitive performance across major multi-modal benchmarks when compared to state-of-the-art specialist-modal LLMs. Specifically, in image benchmarks, Ola achieves 84.3% on MMBench-1.1 [37], 70.8% on MMStar [6], 57.0% on MMMU [73], 68.4% on MathVista [43], 86.1% on AI2D [28] and 827 on OCRBench [38], surpassing all the relative multi-modal LLMs in similar number of parameters. In video benchmarks, Ola attains an impressive 68.4% on VideoMME [20], showcasing its robust capability to handle both video and audio inputs simultaneously, and setting a new state-of-the-art performance among 7B models on the VideoMME benchmark. Ola also maintains a leading position compared to mainstream video LLMs including LLaVA-Video [80] and VideoCCAM [19] on LongVideoBench [66] and MVBench [32]. In audio benchmarks, Ola demonstrates strong audio-speech recognition and conversational abilities, with 3.1% mean WER rate on LibriSpeech [51] and 6.41 mean score on AIR-Bench [70], outperforming existing omni-modal LLMs, including LLaMA-Omni [16], which focuses on audio understanding.
@@ 4.3 Analysis
Analysis on Audio Benchmarks. To demonstrate the effectiveness of our approach on audio and speech tasks, we conducted comprehensive experiments using the LibriSpeech [51], AIR-Bench [70] and MMAU [59] datasets, and we illustrate our results on Tab. 2. Specifically, we report the Word Error Rate (WER) on the test-clean, test-other, dev-clean, and dev-other subsets of LibriSpeech, the GPT-4 eval scores on speech, sound, music, and mix sub-metrics in AIR-Bench, testmini results for MMAU, and GPT-4 eval scores on speech, sound and music for AIR-Foundation. Our model, Ola, is compared against state-of-the-art audio models and omni-modal LLMs.
Notably, Ola demonstrates a significant advantage over existing omni-modal models, achieving a 6.4 average score on AIR-Bench, a 70.3 score on MMAU testmini dataset, and a 60.8 average score on AIR-Foundation. This is in contrast to the previous state-of-the-art omni-modal models, which achieved a 6.4 average score, a 70.3 score, and a 68.8 average score, respectively. Ola’s performance is even approaching that of audio-specific models, highlighting its strong universality.
@@
Figure 5: Generative results on speech and visual understanding tasks. We show the strong ability of omni-modal Ola compared with conventional vision-language models.
@@
Showcases. We present the qualitative generation results of the Ola model on both speech and visual understanding tasks in Fig. 5. For speech understanding, we utilize sources from the AIR-Bench [70] benchmark, where Ola demonstrates precise speech recognition and effective emotional analysis, as well as reasoning about the questions posed. In visual understanding tasks, we analyze an interview with a famous tennis player after the Olympic Games. We compare our approach with the state-of-the-art vision-language model [31]. The large vision-language models without audio inputs exhibit significant information loss due to the absence of audio input and recognition capabilities for audio. In contrast, Ola, with its omni-modal inputs, provides more accurate responses regarding nationality, context, and background from the speaker’s dialogue.
@@ 5 Conclusion
In this paper, we propose Ola, a comprehensive and powerful omni-modal language model that achieves competitive performance in image, video, and audio understanding tasks. Our solution on architectural design, data curation, and training strategy offers a natural, efficient, and competitive pipeline for building an omni-modal model. Enhancements in architectural design with omni-modal inputs, along with high-quality pre-training and fine-tuning data preparation, extend Ola’s capabilities. The training strategy based on joint modality learning and progressive modality alignment connects all the modalities uniformly. We hope our work inspires future research on more general AI models.
@@
We provide supplementary documents to support our research. In the appendix, we provide more details, including model details, training details, and data details in Sec. A. Furthermore, we provide more analysis and more showcases in Sec. B and Sec. C.
@@ A.1 Model Details
The visual encoder for Ola is based on the SigLIP-400M [75] backbone and is further fine-tuned for native-resolution visual inputs. The patch size for the visual encoder is set to 16, the hidden dimension is 1152, and the MLP hidden dimension is 4304. The SigLIP-400M model consists of 27 transformer blocks and 16 attention heads. The audio encoder for Ola is built on the whisper-v3-large [55] and the BEATs [8] model. The length of the input audio tensor for the audio encoder model is fixed at 480,000; therefore, we chunk the entire audio tensor into pieces and concatenate the audio features. The mel size for the whisper-v3-large model is set to 128, and the hidden dimension for the speech features after the whisper-v3-large model is 1280.
@@
We integrate Qwen-2.5-7B [62] into the Ola large language model. The Qwen-2.5-7B model has a hidden LLM dimension of 3,584 and an intermediate size of 18,944. It consists of 28 transformer layers.
@@ A.2 Training Details
As stated in the main paper, the progressive modality alignment procedure is conducted in three stages. The first image-text stage involves adapter pre-training and supervised fine-tuning. The adapter pre-training stage is conducted on 808k image captioning data collected from the LAION datasets. During pre-training, we unfreeze the parameters for the connector while keeping other parameters frozen. We set the training batch size to 256 and the overall learning rate to 1e-3. The supervised fine-tuning stage is conducted on 7.3 million image-text data pairs. During image-text training, input images are maintained at their original aspect ratio, with the maximum image size restricted to 1536×1536. We set the batch size to 128 and the overall learning rate to 2e-5. The stage 1 experiment is conducted on 64 NVIDIA A800 GPUs.